OpenStudy (anonymous):

Expand the series and Evaluate: Enter your answer as the example: 1+2+3+4+5=15 sum_(k=1)^5(k^3)

OpenStudy (mr.math):

\[\sum_{k=1}^5 k^3=1^3+2^3+3^3+4^3+5^3=225.\]

OpenStudy (anonymous):

Thank you!! You are awesome Mr. Math!

OpenStudy (amistre64):

its just a matter of following the directions :)

OpenStudy (mr.math):

@jesusfreak: I know, but thanks anyways :D

OpenStudy (anonymous):

:) @Mr.Math @amistre64, i didn't get it. I don't understand these things

OpenStudy (anonymous):

I have another one though :)